If I'm given a line with a slope of 1/-3 that passes through the point (2, -3)
I need to use point slope form and find the slope intercept form of the line perpendicular to the line at (6,7)
7--3 which becomes 10= 3/1(6-2) but that isn't right I'm so confused please help!

Answer by stanbon(75887)

If I'm given a line with a slope of 1/-3 that passes through the point (2, -3)
I need to use point slope form and find the slope intercept form of the line perpendicular to the line at (6,7)
---------------------------------------
Any line perpendicular to the given line must have slope = 3
because the product of the two slopes must be -1.
---------------------
You want the equation of the line with m= 3 passing thru (6,7)
Use y = mx + b
7 = 3*6 + b
b = 7-18
b = -11
EQUATION:
y = 3x - 11
=================
Cheers,
Stan H.